# Database Views as Explanations for Relational Deep Learning

### v1

@@ 5. GNN-Specific Approach
In this section, we focus on models that are based on hetero-GNNs, as defined in Section 2, and propose a heuristic approach to solving the optimization in Definition 3. One approach to face this challenge is to exhaustively search over the space of candidate explanation views and estimate the deviation from determinacy of each candidate. However, such direct optimization is intractable, as the space of possible views is prohibitively large. As an alternative, we leverage the fact that deep learning libraries are naturally equipped with gradient-based continuous optimization; we use this optimization to efficiently discover locally optimal explanation views via a surrogate continuous optimization problem.
@@ 5.1. Mask Learning: From Discrete to Continuous Optimization
To formulate a continuous optimization problem, we adopt mask learning, a well-established technique for model interpretability (Fong and Vedaldi, 2017; Ying et al., 2019). Mask learning introduces real-valued learnable masks that can then be mapped to candidate explanations in $\mathbf{EL}$, as we will see shortly. In general, mask learning leverages the fact that deep learning models operate with internal feature representations. Masking means selectively hiding some features while retaining others, ideally keeping only the most important ones. [^1] Notice that we use the term masking in the common ML “selector” convention where higher mask values reveal more of the original signal. Concretely, let $\mathord{\mathbf{x}}\in\mathbb{R}^{d}$ be a feature vector. Applying a mask component $m\in[0,1]$ to $\mathord{\mathbf{x}}$ means replacing $\mathord{\mathbf{x}}$ with a new feature vector as follows:
$\mu(\mathord{\mathbf{x}},m)=m\cdot\mathord{\mathbf{x}}+(1-m)\cdot\mathord{\mathbf{u}},$
where $\mathord{\mathbf{u}}$ is a replacement vector of the same dimension as $\mathord{\mathbf{x}}$. If $m=0$, this replaces $\mathord{\mathbf{x}}$ with $\mathord{\mathbf{u}}$. In addition, if $\mathord{\mathbf{u}}$ is uninformative (e.g., the zero vector), masking removes the influence of $\mathord{\mathbf{x}}$. Multiple mask components can be applied in the model simultaneously on different feature vectors $\mathord{\mathbf{x}}$. The collection of these mask components is a mask $\mathord{\mathbf{m}}$. For any value of the mask $\mathord{\mathbf{m}}$, we can obtain a model output, enabling mask components to be learned via gradient descent with an appropriate loss function.
@@ 5.2. Language-Specific Masking
(1) Projection
For projection explanations, we apply mask components attribute-wise in the feature encoding space. Since projection explanations necessarily include all primary and foreign keys, we restrict masks to non-key attributes.
For each relation $R\in\mathcal{S}$, we assign one mask component $m_{R.A_{i}}\in[0,1]$ per attribute $A_{i}\in\mathsf{att}(R)\setminus\big{(}\mathsf{key}(R)\cup\mathsf{FK}(R)\big{)}$. For each node $v=(R,t)$, the same mask vector $(m_{R.A_{1}},\ldots,m_{R.A_{k}})$ is applied at feature encoding, changing Equation 2 as follows:
$\begin{split}\mathord{\mathbf{X}}_{v}=\mathsf{Enc}_{R}(&\mu(\mathsf{Enc}_{R.A_{1}}(t[A_{1}]),m_{R.A_{1}})\oplus\cdots\oplus\\ &\mu(\mathrm{Enc}_{R.A_{k}}(t[A_{k}]),m_{R.A_{k}}))\end{split}$
Recall that $\mu(\mathord{\mathbf{x}},m)$ applies the mask component $m$ to the encoding vector $\mathord{\mathbf{x}}$. When masking attribute $A_{i}$, we use a replacement vector $\mathord{\mathbf{u}}=\mathsf{Enc}_{R.A_{i}}(t^{\prime}[A_{i}])$ for a uniformly and randomly sampled tuple $t^{\prime}\in D(R)$. This ensures the replacement of the attribute’s encoding is uninformative, yet realistic. Masks for all relations and attributes are learned jointly.
After mask learning, continuous mask values are mapped into projection explanations. For that, we use a threshold $\delta>0$:
$m^{\prime}_{R.A_{i}}=\begin{cases}1&\text{if }m_{R.A_{i}}\geq\delta\\ 0&\text{otherwise}\end{cases}$
Attributes with $m^{\prime}_{R.A_{i}}=1$ are included in the projection explanations; the rest are omitted.
(2) FKJoin
For join explanations, we apply masks during message-passing. Every primary/foreign-key pair corresponds to an edge type in the graph (see Section 2.4). Thus, we learn a mask value $m_{\lambda_{i}}\in[0,1]$ for every edge type $\lambda_{i}$ in order to decide whether to retain the corresponding join in the explanation. Equation 1 is adapted as follows:
@@
Notice that for each edge type $\lambda_{i}$, the same mask component $m_{\lambda_{i}}$ is applied at the aggregate vector of messages irrespective of the direction $o_{i}$, as both directions correspond to the same join. Also, the same mask components are used across all layers $\ell$, because message-passing corresponds to the same join operation irrespective of the GNN layer. We replace each aggregate with a zero replacement vector $\mathord{\mathbf{u}}$. Similarly to Projection, continuous mask values are mapped to join explanations after thresholding.
(3) Selection
First, consider the case of a single selection predicate $\phi$. We assign a mask $m_{\phi}\in[0,1]$ to all tuples that satisfy $\phi$. For each node $v=(R,t)$ with $\phi(t)=1$, we set $m_{(R,t)}=m_{\phi}$. Masking is then applied per tuple by adapting Equation 2 into:
$\mathord{\mathbf{X}}_{v}=\mu(\mathsf{Enc}_{R}\left(\mathsf{Enc}_{R.A_{1}}(t[A_{1}])\oplus\cdots\oplus\mathrm{Enc}_{R.A_{k}}(t[A_{k}])\right),m_{(R,t)})$
Similarly to Projection, when masking $v=(R,t)$, we replace $\mathord{\mathbf{X}}_{v}$ with $\mathord{\mathbf{u}}=\mathord{\mathbf{X}}_{v^{\prime}}$ for a randomly sampled $v^{\prime}=(R,t^{\prime})$, where $t^{\prime}\in D(R)$.
For the implementation of Selection, we consider conditions that involve disjunctions of atomic predicates $\phi_{1}\vee\dots\vee\phi_{l}$. Following real-valued logic techniques that map logical formulas into differentiable form (Li et al., 2019), we employ the Łukasiewicz t-conorm for disjunctions. Thus, the mask values per tuple are calculated as $m_{(R,t)}=\min\!\left\{1,\sum_{\phi_{i}(t,\mathord{\mathbf{c}}_{i})=1}m_{\phi_{i}}\right\}$. A similar strategy may be applied for conjunctions. After learning masks for all candidate selection predicates jointly, we apply thresholding to map them into discrete explanations.
To combine selections and projections in a composite language, namely ProjSelect, we first learn projection masks as discussed for Projection, and then learn masks for candidate selection predicates that use data attributes present in the learned projections. In our experiments (Section 6), we restrict ourselves to categorical attributes with equality predicates and numerical attributes with range predicates defined via quantiles.
5.3. Mask Optimization
We define the loss function for mask learning, a surrogate optimization for our original problem in Definition 3:
$\min_{\mathord{\mathbf{m}}}\ \mathcal{L}_{\text{task}}+\lambda\,\mathcal{L}_{\text{reg}}(\mathord{\mathbf{m}}),$
where the mask $\mathord{\mathbf{m}}$ is a vector that contains all mask components to be learned, $\mathcal{L}_{\text{task}}$ is the original loss of the model (e.g., classification or regression loss) and $\mathcal{L}_{\text{reg}}$ is an $\ell_{1}$-regularization term applied to $\mathord{\mathbf{m}}$ that minimizes the size of the mask. Note that $\mathcal{L}_{\text{task}}$ ensures that the learned masks preserve the model’s behavior, a goal that is aligned with the soft determinacy objective of Definition 3. On the other hand, the $\ell_{1}$ penalty encourages sparsity in the masks (Hastie et al., [n.d.]), promoting concise explanations by selecting only the most relevant database components.
@@ 6. Experimental Evaluation
We assess the performance of our proposed method and the overall effectiveness of the framework.
Metrics
@@
The main metric we consider is deviation from determinacy $\mathsf{dev}_{\Delta}(E)$ (see Definition 2). The distance dist is the absolute difference for binary classification and the normalized absolute difference $\textrm{dist}(x,y)=\frac{|x-y|}{|x|+|y|}$ for regression, both bounded in $[0,1]$. Each $\mathsf{dev}_{\Delta}(E)$ estimate is averaged over 5 perturbation samples; we report both mean and standard deviation. Second, we evaluate the explanation size $k$. Recall that this is the number of projected data attributes for Projection, the number of join conditions for FKJoin, and the number of predicates for Selection. Finally, we report the time to generate the explanation, which corresponds to training time for mask-based methods.
Data
@@
We use RelBench, the standard benchmark suite for RDL (Robinson et al., 2024), which includes real-world relational databases and diverse predictive tasks. We focus on node-level tasks, i.e., regression and binary classification. Details of the datasets are provided in the online appendix. As Table 2 shows, we denote each combination of database $i$ and task $j$ as DiiTjj.
Models to explain
@@
We train RDL models from scratch. For each database, we construct a graph representation (as explained in Section 2.4) and use PyTorch Frame (Hu et al., 2024) to implement feature encoders. The GNN is implemented in PyG (Fey et al., 2025), with 32-dimensional encoders. We perform hyperparameter tuning over the learning rate, batch size, and number of GNN layers. Models are trained using the train/validation/test split provided by the benchmark. Our models reach performance similar to the one previously reported (Robinson et al., 2024) for GNN approaches. For each explanation task, we sample 100 instances for mask training and 100 instances for explanation evaluation. Since many classification tasks are imbalanced, we use balanced sampling to ensure that the selected instances contain an equal number from each class. For temporal tasks we perform temporal sampling (Fey et al., 2024) to ensure time-consistency.
Methods
1
@@ Methods
Mask-based: We implement the learnable masks as introduced in Section 5, with early stopping. 1a Column Mask: masks for Projection. 1b FKPK Mask: masks for FKJoin. 1c Filter Mask: masks for Selection.
@@
Since no prior work generates SQL explanations for RDL, we implement a set of baselines. 2 Ranking-based for Projection: We implement two baselines that order data attributes according to their importance, and select the top-$k$ for each target explanation size $k$. A known limitation of such methods is that they ignore dependencies between attributes (Molnar, 2025). 2a Local Impact: The importance of a data attribute is computed as $dev_{\Delta}$ when the attribute is taken as an explanation by itself. The lower the $dev_{\Delta}$ the more important the attribute. 2b PFI: We adapt Permutation Feature Importance (PFI) (Molnar, 2025) by measuring the change in $dev_{\Delta}$ when each attribute is removed. Attributes are ranked by impact, where higher deviation implies higher importance.
3
@@
Greedy approach for Projection and FKJoin: 3a Greedy: Starting from the empty set, we iteratively add the attribute that yields the largest reduction in $dev_{\Delta}$. 3b Greedy Expansion: At each step, we add a join that maximally reduces $dev_{\Delta}$ while ensuring connectivity to the prediction entity. This is more computationally feasible than Greedydue to fewer join candidates at each iteration.
5
@@
Random Subset for Projection: We randomly sample the data attributes and report the average over 5 samples.
4
Empty: An explanation containing no information, resulting in a full database perturbation. We use this as a point of reference.
Implementation details
@@
For the embeddings of textual attributes we use a BERT-based model [^2] https://huggingface.co/sentence-transformers/distilbert-base-nli-mean-tokens. All experiments are conducted on a VM with an RTX 6000 Ada GPU, 14 vCPUs, and 188 GB of RAM. Our code is available at https://github.com/agapiR/rdl-explain.
Evaluation plan
@@
We evaluate the proposed mask learning strategy separately for the Projection and FKJoin languages. Our goal is to evaluate the tradeoff between conciseness and soft determinacy. To ensure a fair comparison, we use the same explanation size for all methods, denoted as $k^{*}$, which we determine by thresholding the learned mask as discussed in Section 5. We also assess the robustness of $dev_{\Delta}$ under different database perturbation strategies $\Delta$. Additionally, we conduct a case study on one task where we showcase Selection and $3$ composite languages.

### v2

@@ 5. GNN-Specific Approach
We now instantiate our framework (Section 3) for the case of GNNs. We present a model-specific approach that leverages properties of the model to efficiently solve the optimization in Definition 3: differentiability enables explanation discovery via gradient-based learning, while the model structure (i.e., feature encoding followed by message-passing) provides natural intervention points.
@@ 5.1. Mask Learning: From Discrete to Continuous Optimization
As discussed in Section 3.3, a direct solution to Definition 3 is not practically feasible. We now show how mask learning naturally arises as a continuous relaxation of the discrete soft determinacy problem, focusing on Projection explanations.
Discrete Selection Problem. Recall that soft determinacy (Definition 2) measures sufficiency of an explanation $E$ by evaluating the expected distance between $M_{D}(s)$ and $M_{D^{\prime}}(s)$ over contingency databases $D^{\prime}$ that respect the explanation views. For Projection explanations over $N$ data attributes, we seek a binary selection vector $\mathord{\mathbf{a}}\in\{0,1\}^{N}$ indicating which attributes to include in the explanation. To evaluate soft determinacy, we construct contingency databases by applying the joint column-wise permutation: for each tuple $t$, we replace the value of each attribute $A_{i}$ as:
$t[A_{i}]\leftarrow a_{i}\cdot t[A_{i}]+(1-a_{i})\cdot t^{\prime}[A_{i}]$
where $t^{\prime}$ is a randomly selected tuple from the same relation, and $a_{i}\in\{0,1\}$ determines whether attribute $A_{i}$ is fixed ($a_{i}=1$) or replaced ($a_{i}=0$). The soft determinacy objective then becomes:
[…]
where $s$ is an instance, $\omega$ represents the randomness in selecting replacement tuples $t^{\prime}$, and $\mathord{\mathbf{a}}||_{0}$ counts the number of selected attributes. However, optimizing over $\mathord{\mathbf{a}}\in\{0,1\}^{N}$ requires searching through $2^{N}$ combinations, which is intractable.
Continuous Relaxation via Direct Masking. A natural approach is to relax the binary constraint: replace $\mathord{\mathbf{a}}\in\{0,1\}^{N}$ with continuous mask values $\mathord{\mathbf{m}}\in[0,1]^{N}$. This yields the continuous optimization problem:
[…]
where we apply the masking operation at the feature encoding level. Concretely, let $\mathord{\mathbf{x}}\in\mathbb{R}^{d}$ be a feature vector for attribute $A_{i}$. Applying mask component $m_{i}\in[0,1]$ means replacing it with:
$\mu_{\omega}(\mathord{\mathbf{x}},m_{i})=m_{i}\cdot\mathord{\mathbf{x}}+(1-m_{i})\cdot\mathord{\mathbf{u}}_{\omega},$
where $\mathord{\mathbf{u}}_{\omega}$ is the encoding of the replacement value from tuple $t^{\prime}$, i.e., $\mathord{\mathbf{u}}_{\omega}=\mathsf{Enc}_{T.A_{i}}(t^{\prime}[A_{i}])$. Note that the $\ell_{1}$ penalty $\mathord{\mathbf{m}}||_{1}$ serves as a convex relaxation of the $\ell_{0}$ cardinality constraint, encouraging sparsity (Hastie et al., 2015). Additionally, the function $\text{loss}()$ is now a continuous distance metric (e.g., cross entropy, mean squared error).
This continuous relaxation enables gradient-based optimization using automatic differentiation frameworks. When $m_{i}=0$, attribute $A_{i}$ is completely replaced, and as $m_{i}$ increases toward 1, the original value is progressively restored. In practice, we use stochastic optimization (Kingma and Ba, 2015) as follows: for each optimization step we first sample a mini-batch of instances from $D(R)$ to compute the empirical loss, and then we apply one realization of the random replacement $\omega$ (e.g., one random permutation of tuples in each relation of $D$). After optimization, we threshold masks using $m_{i}\geq\delta$ to recover discrete explanations, a common technique for mask discretization (Amara et al., 2022).
Alternative Approaches. While direct masking has proven effective for interpretability in images (Fong and Vedaldi, 2017) and graph neural networks (Ying et al., 2019), other techniques also exist for continuous relaxation of discrete selection problems. Policy gradient methods such as REINFORCE (Williams, 1992) treat selection as a stochastic policy, computing gradients via the likelihood ratio trick. Reparameterization approaches like Gumbel-Softmax (Jang et al., 2017; Maddison et al., 2014, 2016) introduce auxiliary random variables to create differentiable approximations of discrete sampling. For example, L2X (Chen et al., 2018) uses these techniques to compute local explanations via feature selection with a fixed budget.
We adopt direct masking due to several practical advantages. First, it provides computational efficiency: requiring only a single forward pass per mask-learning step without sampling auxiliary variables, and producing stable gradients at all points in the continuous domain $[0,1]$. Second, it naturally fits for the complex RDL model, where computation involves multiple stages (feature encoding, multi-layer message passing), allowing unified intervention across different explanation languages $\mathbf{EL}$ by modulating continuous internal representations. The latter enables joint optimization for composite languages, as the same masking operation applies regardless of whether we mask attributes (Projection), foreign keys (FKJoin), or tuple groups (Selection).
Quality of Mask Learning Solutions. We discussed how mask learning arises as a natural continuous relaxation of the discrete soft determinacy problem. Specifically, for Projection and discrete masks ($\mathord{\mathbf{m}}\in\{0,1\}^{n}$), the objectives are equivalent: $\mathsf{dev}_{\Delta^{\mathsf{joint}}_{E}}(E)=\mathbb{E}_{s}\,\mathbb{E}_{\omega}\left[\text{loss}(M_{D}(s),M_{\mathord{\mathbf{m}},\omega,D}(s))\right]$. In practice, we learn continuous masks and threshold to obtain discrete explanations. We now argue that this approach yields high-quality solutions.
Consider the simplified setting of a binary classifier $M_{\mathcal{C}}$ that relies on precisely $k$ data attributes. First, there exists a high-quality Projection explanation $E^{*}$ for $M_{\mathcal{C}}$: since $M_{\mathcal{C}}$ uses exactly $k$ columns, the $k$-column explanation selecting these achieves cost $\ell^{*}=\lambda\cdot k$ (zero task loss plus regularization). Second, the mask learning objective promotes this solution: any smaller mask (with $k^{\prime}<k$ attributes) will incur substantial task loss since important features are masked, thus predictions flip with probability $\epsilon>0$, yielding expected cost $\ell^{\prime}=\epsilon+\lambda\cdot k^{\prime}>>\ell^{*}$ for appropriate $\lambda$. Conversely, any larger mask (with $k^{\prime\prime}>k$ attributes) incurs higher regularization cost $\lambda\cdot k^{\prime\prime}>\ell^{*}$ without reducing task loss. Hence, the mask learning objective promotes the optimal $k$-column explanation $E^{*}$.
@@ 5.2. Language-Specific Masking
(1) Projection. We apply mask components attribute-wise in the feature encoding space. Since Projection necessarily includes all primary and foreign keys, we restrict masks to data attributes.
For each relation $R\in\mathcal{S}$, we assign one mask component $m_{R.A_{i}}\in[0,1]$ per data attribute $A_{i}$, $i=1,\dots,n$. For each node $v=(R,t)$, the same mask vector $(m_{R.A_{1}},\ldots,m_{R.A_{n}})$ is applied at feature encoding, changing Equation 2 as follows:
$\begin{split}\mathord{\mathbf{X}}_{v}=\mathsf{Enc}_{R}(&\mu(\mathsf{Enc}_{R.A_{1}}(t[A_{1}]),m_{R.A_{1}})\oplus\cdots\oplus\\ &\mu(\mathrm{Enc}_{R.A_{n}}(t[A_{n}]),m_{R.A_{n}}))\end{split}$
Recall that $\mu(\mathord{\mathbf{x}},m)$ applies the mask component $m$ to the encoding vector $\mathord{\mathbf{x}}$. Every time we apply masking attribute-wise, we first consider a random permutation to the tuples of $D(R)$, which assigns for each tuple $t$ a replacement tuple $t^{\prime}$. Then to apply masking to $v=(R,t)$ for each attribute $A_{i}$ we use the replacement vector $\mathord{\mathbf{u}}_{i}=\mathsf{Enc}_{R.A_{i}}(t^{\prime}[A_{i}])$, i.e., the encoding of the value of $A_{i}$ in $t^{\prime}$. This ensures the replacement of the $t[A_{i}]$ encoding is uninformative, yet realistic. Masks for all relations and attributes are learned jointly.
(2) FKJoin. We apply masks at message-passing, learning a mask value $m_{\lambda_{i}}\in[0,1]$ for every edge type $\lambda_{i}$. Since every foreign-key pair corresponds to an edge type (see Section 2.4), $m_{\lambda_{i}}$ indicates the importance of the corresponding join. Equation 1 becomes:
@@
For each edge type $\lambda_{i}$, the same mask component $m_{\lambda_{i}}$ is applied at the aggregate vector of messages irrespective of the direction $o_{i}$, as both directions correspond to the same join. Also, the same mask components are used across all layers $\ell$, because message-passing corresponds to the same join operation irrespective of the GNN layer. We replace each aggregate with a zero replacement vector $\mathord{\mathbf{u}}$.
(3) Selection. First, consider the case of a single selection predicate $\phi$. We assign a mask $m_{\phi}\in[0,1]$ to all tuples that satisfy $\phi$. For each node $v=(R,t)$ with $\phi(t)=1$, we set $m_{(R,t)}=m_{\phi}$. Masking is then applied per tuple by adapting Equation 2 into:
$\mathord{\mathbf{X}}_{v}=\mu(\mathsf{Enc}_{R}\left(\mathsf{Enc}_{R.A_{1}}(t[A_{1}])\oplus\cdots\oplus\mathrm{Enc}_{R.A_{n}}(t[A_{n}])\right),m_{(R,t)})$
Similarly to Projection, when masking $v=(R,t)$, we replace $\mathord{\mathbf{X}}_{v}$ with $\mathord{\mathbf{u}}=\mathord{\mathbf{X}}_{v^{\prime}}$ for a replacement $v^{\prime}=(R,t^{\prime})$, $t^{\prime}\in D(R)$ according to a random permutation of tuples in $D(R)$.
For the implementation of Selection, we consider conditions that involve disjunctions of atomic predicates $\phi_{1}\vee\dots\vee\phi_{l}$. Following real-valued logic techniques that map logical formulas into differentiable form (Li et al., 2019), we employ the Łukasiewicz t-conorm for disjunctions. Thus, the mask values per tuple are calculated as $m_{(R,t)}=\min\!\left\{1,\sum_{\phi_{i}(t,\mathord{\mathbf{c}}_{i})=1}m_{\phi_{i}}\right\}$. A similar strategy may be applied for conjunctions. In our experiments (Section 6), we restrict ourselves to categorical attributes with equality predicates and numerical attributes with range predicates defined via quantiles.
In all cases, continuous mask values are mapped into discrete explanations using a threshold $\delta>0$; $m^{\prime}_{x}=1$ if $m_{x}\geq\delta$ and 0 otherwise. Components $x$ with $m^{\prime}_{x}=1$ are included in the explanations; the rest are omitted.
@@ 6. Experimental Evaluation
Metrics. The main metric is deviation from determinacy $\mathsf{dev}_{\Delta}(E)$ (see Definition 2). The distance dist is absolute difference for binary classification and normalized absolute difference $\textrm{dist}(x,y)=\frac{|x-y|}{|x|+|y|}$ for regression, both bounded in $[0,1]$. Each $\mathsf{dev}_{\Delta}(E)$ estimate is over 5 perturbation samples; we report both mean and standard deviation. Second, we evaluate the explanation size $k$. Recall that this is the number of projected data attributes for Projection, the number of join conditions for FKJoin, and the number of predicates for Selection. Finally, we report the time to generate the explanation, which corresponds to training time for mask learning.
@@
Data. We use RelBench, the standard benchmark for RDL (Robinson et al., 2024) with real-world databases and diverse predictive tasks. We focus on node-level tasks, i.e., regression and binary classification. Dataset details are provided in the online appendix. As Table 2 shows, we denote each combination of database $i$ and task $j$ as DiiTjj.
@@
Explained Models. We train RDL models from scratch. For each database, we construct a graph (Section 2.4) and use PyTorch Frame (Hu et al., 2024) for feature encoders. The GNN is implemented in PyG (Fey et al., 2025) with 32-dimensional channels. We perform hyperparameter tuning over the learning rate, batch size, and number of GNN layers, and use the train/val/test split of the benchmark. Our models reach performance similar to the one previously reported (Robinson et al., 2024) for GNN approaches. For each explanation task, we sample 100 instances for mask training and 100 instances for evaluation. For classification tasks, we report a class-balanced $dev_{\Delta}$ that averages over an equal number of instances from each class, avoiding majority-class dominance. For temporal tasks we perform temporal sampling (Fey et al., 2024) to ensure time-consistency.
@@
Methods. (1) Mask-based: Our method as introduced in Section 5, with early stopping and thresholding with $\delta=0.1$. Note that, on average, only 1.97% of mask values for Projection and 2.17% for FKJoin fall in the range $[0.05,0.2]$ and are sensitive to small $\delta$ variations. (1a) Column Mask: Masks for Projection. (1b) FKPK Mask: Masks for FKJoin. (1c) Filter Mask: Masks for Selection.
@@
Since no prior work generates SQL explanations for RDL, we implement a set of baselines. (2) Ranking-based for Projection: We order data attributes according to their importance, and select the top-$k$ for each target explanation size $k$. A known limitation of such methods is that they ignore dependencies between attributes (Molnar, 2025). (2a) Local Impact: The importance of a data attribute is computed as $dev_{\Delta}$ when the attribute is taken as an explanation by itself. The lower the $dev_{\Delta}$ the more important the attribute. (2b) PFI: We adapt Permutation Feature Importance (PFI) (Molnar, 2025) by measuring the change in $dev_{\Delta}$ when each attribute is removed. Attributes are ranked by impact, where higher deviation implies higher importance.
@@
(3) Greedy approach for Projection and FKJoin: (3a) Greedy: Starting from the empty set, we iteratively add the attribute that yields the largest reduction in $dev_{\Delta}$. (3b) Greedy Expansion: At each step, we add a join that maximally reduces $dev_{\Delta}$ while ensuring connectivity to the prediction entity. This is more computationally feasible than Greedy due to fewer join candidates at each iteration.
@@
(4) Random Subset for Projection: We randomly sample the data attributes and report the average over 5 samples.
@@
Experiment details. For text attribute embeddings we use a BERT-based model [^1] https://huggingface.co/sentence-transformers/distilbert-base-nli-mean-tokens. Experiments are done on a VM with an RTX 6000 Ada GPU, 14 vCPUs, and 188 GB of RAM. Our code is public [^2] https://github.com/agapiR/rdl-explain.
@@
Evaluation plan. We evaluate mask learning separately for Projection and FKJoin. For fair comparison, we use the same explanation size for all methods, denoted as $k^{*}$, which we determine by mask thresholding as discussed in Section 5. We also assess the robustness of $dev_{\Delta}$ under different database perturbation strategies $\Delta$. Finally, we conduct a case study to show end-user usefulness.